# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

### v5

@@ Abstract
Offering a promising solution to the scalability challenges associated with human evaluation, the LLM-as-a-judge paradigm is rapidly gaining traction as an approach to evaluating large language models (LLMs). However, there are still many open questions about the strengths and weaknesses of this paradigm, and what potential biases it may hold. In this paper, we present a comprehensive study of the performance of various LLMs acting as judges [^1] Source code is available at https://github.com/UMass-Meta-LLM-Eval/llm_eval Investigating thirteen judge models of different model sizes and families, judging answers of nine different ‘exam-taker models’ – both base and instruction-tuned – we find that only the best (and largest) models achieve reasonable alignment with humans. However, they are still quite far behind inter-human agreement and their assigned scores may still differ with up to 5 points from human-assigned scores. In terms of their ranking of the nine exam-taker models, instead, also smaller models and even the lexical metric contains may provide a reasonable signal. Through error analysis and other studies, we identify vulnerabilities in judge models, such as their sensitivity to prompt complexity and length, and a tendency toward leniency. The fact that even the best judges differ from humans in this comparatively simple setup suggest that caution may be wise when using judges in more complex setups. Lastly, our research rediscovers the importance of using alignment metrics beyond simple percent alignment, showing that judges with high percent agreement can still assign vastly different scores.
@@ 1 Introduction
Figure 1: Average scores assigned by judge models and alignment with human judges. (a) Scores assigned to all exam-taker models by the various judge models. (b) Average percent agreement (blue line) and Scott’s pi scores (red bars) of judge models with human judges (black line). Error bars annotate standard deviation across exam-taker models. Llama3 70B, Llama3.1 70B and GPT-4 Turbo have Scott’s pi coefficient that are indicative of excellent alignment, but are still well below the human alignment score.
@@
Over the last few years, large language models (LLMs) have demonstrated remarkable capabilities across various domains (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; AI@Meta, 2024, i.a.). As more and more new LLMs with different architectures and training methods continue to be released and their capabilities expand, accurately evaluating their performance and limitations becomes increasingly challenging (Zheng et al., 2024; Ohmer et al., 2024; Benchekroun et al., 2023; Madaan et al., 2024). The empirical evaluation of LLMs is particularly difficult due to the diversity of their outputs and the wide range of tasks they are used for (Zhang et al., 2024; Li et al., 2023a).
To evaluate LLMs, various methods have been proposed, typically falling into one of two broad categories. First, benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), or GSM8K (Cobbe et al., 2021) are used to evaluate specific capabilities of LLMs in an automated manner. Additionally, leaderboards like Chatbot Arena (Chiang et al., 2024) and Open LLM Leaderboard (Beeching et al., 2023) assign ranks to models considering pair-wise rankings of LLM outputs, done by humans or, in some cases, automated evaluation methods. Since both strategies involve evaluating free-form text responses generated by the LLMs, even in the first case, evaluating the responses is often just as challenging as generating them (see e.g. Chang et al., 2023; Bavaresco et al., 2024).
@@
One proposed solution to this problem is to use multiple-choice question (MCQ) benchmarks such as MMLU, and compare the log-probabilities of the potential answers rather than evaluating the generated answer directly. However, the MCQ paradigm limits the range of abilities that can be evaluated, and the setup increasingly diverges from how LLMs are used in practice. Alternatively, the use of lexical matching methods such as exact match (EM) or n-gram overlap to evaluate the responses are practical and cost-efficient approaches, but are susceptible to false negatives and often fail to adequately distinguish between responses with subtle differences that change their semantic meaning. This issue is exacerbated when evaluating instruction-tuned “chat” models that are fine-tuned to carry out conversations with humans in natural language, since their responses tend to be more verbose (Saito et al., 2023; Renze & Guven, 2024). For these reasons, human evaluation remains the gold standard for evaluating LLM responses.
@@
However, human evaluation is expensive, time-consuming, and often impractical in many use cases. As a result, it has increasingly become common practice to evaluate LLM responses using another LLM as a judge model (Lin et al., 2021; Islam et al., 2023; Chiang & Lee, 2023; Liusie et al., 2024). While there are promises of alignment between LLM judges and humans (Sottana et al., 2023; Zheng et al., 2024), there are also many open questions about the strengths and weaknesses of the paradigm. In this work, we study the properties of LLMs as judges, comparing them with humans and automated evaluation methods. Contrary to prior work, we focus on a clean scenario in which human alignment is very high, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models. Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how thirteen different judge models with varying architectures and sizes judge nine different exam-taker models. Our main findings are:
@@
  • Even in clean and straightforward setups, only the best models have high alignment scores. Out of the thirteen judge models we considered, only GPT-4 Turbo, Llama-3.1 70B and Llama-3 70B showed very high alignment with humans. Also for those judges, though, alignment is still well behind the human alignment coefficient for the task (Figure 1).
@@
  • Scott’s $\mathbf{\pi}$ distinguishes judges better than percent alignment. In terms of percent alignment, judges are rarely discriminable, while Scott’s $\mathbf{\pi}$ provides a more informative signal. In some cases, high percent agreement can still give scores that differ 10-20 points from the human-assigned scores (Figure 2).
@@
  • Also Scott’s $\mathbf{\pi}$ is not all telling. While GPT-4 Turbo and Llama-3 both have alignment scores that are considered excellent, their scores still differ up to 5 points from human-assigned scores. Furthermore, when it comes to discriminating different exam-taker models, their results are comparable to alternative cheaper approaches such as Mistral 7B and contains, which have much lower alignment scores but more consistent biases (Figure 3).
@@
Through detailed analysis (§ 5), we uncover additional insights into judge performance. Improved alignment appears to be driven by improved recall rates and reduced false negatives. However, judge models struggle with under-specified answers and tend to be lenient, affecting their evaluation consistency. They are also sensitive to the length and quality of prompts. And, surprisingly, even when the judge models are asked to evaluate an answer matching verbatim with a reference answer, many judge models still sometimes fail to evaluate it correctly.
@@
Overall, our work showcases the strengths of the LLM-as-a-judge paradigm while also highlighting the need for caution against overreliance on alignment metrics, even in cases where they are high. Through error analysis, we also highlight several common failure cases that require attention. With this, we aim to contribute to a better general understanding of what is now becoming a mainstream paradigm for evaluating LLMs.
@@ 2 Related work
Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang & Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024). Zhang et al. (2024) and Sottana et al. (2023) propose ways to standardise LLM evaluations and the role that judge models might play in such solutions. Several studies have demonstrated that state-of-the-art LLMs such as GPT-4 Turbo exhibit high alignment with human judgments (Sottana et al., 2023; Zheng et al., 2024), though others also illustrate that the paradigm is not yet without faults. Zeng et al. (2023) propose a benchmark for evaluating the performance of LLMs as judges, and other approaches have been proposed to improve LLM judges such that they are aligned well with humans (Shankar et al., 2024; Zhu et al., 2023).
Despite promising results in various settings, judge models still suffer from known issues of current LLMs such as hallucinations and factual errors (Ye et al., 2023; Turpin et al., 2023) and difficulty in following complex instructions (Li et al., 2023b; He et al., 2024). Furthermore, various studies have reported challenges such as position bias (Pezeshkpour & Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu & Aji, 2023). Recently, Liusie et al. (2024) have shown that LLMs perform better in comparative assessment compared to absolute scoring, which can be used for reliably measuring the relative performance of models (Liu et al., 2024) and creating classifiers for pairwise grading (Huang et al., 2024).
We follow up on this line of work and investigate the strengths and weaknesses of LLMs as judges. Unlike most prior work, we do not focus on pairwise comparisons of LLM outputs on open-ended tasks, but on comparisons of LLM outputs and reference answers. Since human alignment is high in this setting, this provides a clean playground to study the strengths and weaknesses of LLMs in detail. We also extend previous work by considering more LLMs, both as judges and LLMs to be evaluated.
Table 1: Exam-taker models and judge models We consider a wide variety of exam-taker models and judge models; to get a in-depth overview of their abilities, we consider exam-taker models of various sizes & types.
Exam-taker models       (base & instruction-tuned) Llama-2 (7B, 13B and 70 B), Mistral 7B, GPT-4 Turbo
Judge models (instruction-tuned) Llama-2 (7B, 13B, 70B),
Llama-3 (8B, 70B), Llama-3.1 (8B, 70B),
Gemma 2B, Mistral 7B, JudgeLM 7B, GPT-4 Turbo
Judge models (lexical) Exact Match (EM), Contains
@@ 3 Methodology
As our testbed, we use the TriviaQA dataset (Joshi et al., 2017), consisting of 95K question-answer pairs sourced from 14 trivia and quiz league websites. Each question in the train and validation set is annotated with a list of short answers containing a minimal set of facts and evidence documents collected from Wikipedia and the Web. For our experiments, we use the validation set of the unfiltered partition of the benchmark, using the short answers as reference answers. We use the training set for few-shot examples. Since experiments require manual annotation of the exam-taker model responses, we use a random sample of $400$ questions from the dataset. In Appendix I, we show with a bootstrapping test that this sample size has low variance for our main result. Through experiments described in § 3, we establish that humans have high agreement on judgements of answers given to the questions in the benchmark.
@@
To understand the strengths and weaknesses of different judges, we consider answers of pre-trained (base) and instruction-tuned (chat) ‘exam-taker models’ across a wide variety of model sizes. In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^2] Accessed via the OpenAI API between Mar 19th, 2024 and Sep 20, 2024. (Achiam et al., 2023) as the exam-taker models. The prompts for the exam-taker models contain five few-shot examples of (question, answer) pairs from the TriviaQA training set. The prompts for the instruction-tuned models additionally include a command signaling the model to answer the given question in a succinct manner similar to the provided examples. The prompts are provided in Appendix D.
@@
To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Llama-3.1 (Dubey et al., 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges. To maintain parity with human and judge evaluation, judge prompts were built from human guidelines in Appendix G. The judges are instructed to respond with only a single word, ‘‘correct’’ or ‘‘incorrect’’. An overview of all exam-taker models and judge models is shown in Table 1. For ease of reading, the judge models are depicted in a different font than the exam-taker models.
@@
We use two metrics to quantify alignment between judges: percent agreement and Scott’s Pi coefficient (Scott, 1955). [^3] In an earlier version of this paper, we used Cohen’s kappa (Cohen, 1960) to measure alignment. It has since come to our attention that – despite it’s widespread use – this metric has some well-documented theoretical issues (e.g. Pontius & Millones, 2011; Chicco et al., 2021). For the interested reader, we elaborate on these issues in Appendix B. Percent agreement expresses a simple percentage of the samples on which two annotators agree. Scott’s Pi, denoted as Scott’s $\mathbf{\pi}$, is an alignment metric that corrects for chance agreement between two annotators and is considered to provide a more robust measure of alignment. Details about the computation of both metrics are given in Appendix F.
@@
As a ground-truth assessment, we obtain human annotations for each exam-taker model answer. The inter-human alignment is calculated between three human judges using the answers to 1200 randomly sampled questions answers; the human guidelines can be found in Appendix G. We then determine collective “Human Judgment” through a majority vote. The average alignment among human evaluators with the majority vote had a Scott’s $\mathbf{\pi}$ of $96.2\pm 1.07$,444coefficient scaled by $100$ for easier comparison with percent alignment. and the average percent agreement was $98.52\%\pm 0.42\%$. The details of this experiment are mentioned in Appendix A. Given this near-perfect alignment score, we consider only one human evaluator per sample for the rest of our experiments, to reduce the overall cost of human annotations. The set of questions for which we obtain human annotations is identical for each exam-taker model.

### v6

@@ Abstract
The LLM-as-a-judge paradigm offers a potential solution to scalability issues in human evaluation of large language models (LLMs), but there are still many open questions about its strengths, weaknesses, and potential biases. This study investigates thirteen models, ranging in size and family, as ‘judge models’ evaluating answers from nine base and instruction-tuned ‘exam-taker models’. We find that only the best (and largest) models show reasonable alignment with humans, though they still differ with up to 5 points from human-assigned scores. Our research highlights the need for alignment metrics beyond percent agreement, as judges with high agreement can still assign vastly different scores. We also find that smaller models and the lexical metric contains can provide a reasonable signal in ranking the exam-taker models. Further error analysis reveals vulnerabilities in judge models, such as sensitivity to prompt complexity and a bias toward leniency. Our findings show that even the best judge models differ from humans in this fairly sterile setup, indicating that caution is warranted when applying judge models in more complex scenarios.
Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges
Aman Singh Thakurthanks: Equal Contributionand Kartik Choudharyfootnotemark: and Venkat Srinik Ramayapallyfootnotemark: University of Massachusetts Amherst {amansinghtha, kartikchoudh, vramayapally}@umass.edu
Sankaran Vaidyanathan University of Massachusetts Amherst sankaranv@cs.umass.edu          Dieuwke Hupkes Meta dieuwkehupkes@meta.com
@@ 1 Introduction
Figure 1: Average scores assigned by judge models and alignment with human judges. (a) Scores assigned to all exam-taker models by the various judge models. (b) Average percent agreement (blue line) and Scott’s $\pi$ scores (red bars) of judge models with human judges (black line). Error bars annotate standard deviation across exam-taker models. Llama3 70B, Llama3.1 70B and GPT-4 Turbo have Scott’s $\pi$ coefficient that are indicative of excellent alignment, but are still well below the human alignment score.
@@
Over the last few years, large language models (LLMs) have demonstrated remarkable capabilities across various domains (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; AI@Meta, 2024, i.a.). As more and more new LLMs with different architectures and training methods continue to be released and their capabilities expand, accurately evaluating their performance and limitations becomes increasingly challenging (Zheng et al., 2024; Ohmer et al., 2024; Benchekroun et al., 2023; Madaan et al., 2024; Li et al., 2023a).
LLM evaluation methods generally fall into one of two broad categories. Benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), and GSM8K (Cobbe et al., 2021) assess specific capabilities, while leaderboards such as Chatbot Arena (Chiang et al., 2024) and Open LLM Leaderboard (Beeching et al., 2023) rank models based on human or automated pairwise comparisons. Both approaches face challenges in evaluating free-form text responses, as assessment can be as difficult as generation itself (see e.g. Chang et al., 2023; Bavaresco et al., 2024).
@@
One approach to evaluating LLMs is using MCQ benchmarks like MMLU, which compare answer log-probabilities instead of assessing generated responses directly. However, this approach limits the range of measurable abilities and differs from how LLMs are used in practice. Lexical methods, such as exact match (EM) or n-gram overlap, are practical and cost-effective but prone to false negatives and often miss subtle semantic differences. These challenges are amplified for instruction-tuned chat models, which tend to produce more verbose responses (Saito et al., 2023; Renze and Guven, 2024).
For these reasons, human evaluation remains the gold standard for evaluating LLM responses.
Exam-taker models (base & instruction-tuned) Llama-2 (7B, 13B, 70B), Mistral 7B, GPT-4 Turbo
Judge models (instruction-tuned) Llama-2 (7B, 13B, 70B), Llama-3 (8B, 70B), Llama-3.1 (8B, 70B), Gemma 2B, Mistral 7B, JudgeLM 7B, GPT-4 Turbo
Judge models (lexical) Exact Match (EM), Contains
Table 1: Exam-taker models and judge models We consider a wide variety of exam-taker models and judge models; to get an in-depth overview of their abilities, we consider exam-taker models of various sizes & types.
@@
Human evaluation is, however, expensive and often impractical, leading to the growing use of LLMs as judge models (Lin et al., 2021; Islam et al., 2023; Chiang and Lee, 2023; Liusie et al., 2024). While promising alignment with humans has been noted (Sottana et al., 2023; Zheng et al., 2024), questions about this approach remain. This work examines LLMs as judges, contrasting them with humans and automated methods. Unlike prior studies, we focus on scenarios with high human alignment to separate task ambiguity from judge model limitations. Using TriviaQA (Joshi et al., 2017), we evaluate how judge models of varying architectures and sizes assess exam-taker models.
In this work, we study the properties of LLMs as judges, comparing them with humans and automated evaluation methods. Contrary to prior work, we focus on a clean scenario in which human alignment is very high, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models. Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how thirteen different judge models with varying architectures and sizes judge nine different exam-taker models. Our main findings are:
@@
  • Even in clean setups, only the best models have high alignment scores. Among the thirteen judge models, only GPT-4 Turbo, Llama-3.1;70B, and Llama-3;70B achieved strong alignment with humans. However, even these fall short of the human alignment coefficient (Figure˜1).
@@
  • Scott’s $\mathbf{\pi}$ distinguishes judges better than percent alignment. In terms of percent alignment, judges are rarely discriminable, while Scott’s $\mathbf{\pi}$ provides a more informative signal. In some cases, high percent agreement can still give scores that differ 10-20 points from the human-assigned scores (Figure˜2).
@@
  • Also Scott’s $\mathbf{\pi}$ is not all telling While GPT-4 Turbo and Llama-3 achieve excellent alignment scores, they can differ by up to 5 points from human scores. Moreover, in discriminating between exam-taker models, their performance is comparable to cheaper alternatives like Mistral 7B and contains, which have lower alignment scores but more consistent biases (Figure˜3).
@@
Through detailed analysis (§˜5), we gain insights into judge performance. Improved alignment appear to be driven from higher recall rates and fewer false negatives. However, judge models struggle with under-specified answers and exhibit leniency, reducing evaluation consistency. They are also sensitive to prompt length and quality. Surprisingly, even when asked to evaluate a verbatim match with a reference, judge models sometimes fail.
@@
Overall, our work highlights the strengths of the LLM-as-a-judge paradigm, while cautioning against overreliance on alignment metrics, even when they are high. Through error analysis, we identify common failure cases, contributing to a deeper understanding of this emerging evaluation paradigm. With this work, our objective is to improve understanding of the emerging mainstream paradigm for evaluating LLM.
@@ 2 Related work
Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang and Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024). Zhang et al. (2024) and Sottana et al. (2023) propose ways to standardise LLM evaluations and the role that judge models might play in such solutions. Several studies have demonstrated that state-of-the-art LLMs such as GPT-4 Turbo exhibit high alignment with human judgments (Sottana et al., 2023; Zheng et al., 2024), though others also illustrate that the paradigm is not yet without faults. Zeng et al. (2023) propose a benchmark for evaluating the performance of LLMs as judges, and other approaches have been proposed to improve LLM judges such that they are aligned well with humans (Shankar et al., 2024; Zhu et al., 2023).
Despite promising results in various settings, judge models still suffer from known issues of current LLMs such as hallucinations and factual errors (Ye et al., 2023; Turpin et al., 2023) and difficulty in following complex instructions (Li et al., 2023b; He et al., 2024). Furthermore, various studies have reported challenges such as position bias (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu and Aji, 2023). Recently, Liusie et al. (2024) have shown that LLMs perform better in comparative assessment compared to absolute scoring, which can be used for reliably measuring the relative performance of models (Liu et al., 2024) and creating classifiers for pairwise grading (Huang et al., 2024).
We build on previous work to investigate the strengths and weaknesses of LLMs as judges. Unlike previous studies, we focus on comparing LLM outputs with reference answers rather than pairwise comparisons on open-ended tasks. With high human alignment in this setting, we gain a clearer view of LLM performance. Furthermore, we extend previous research by considering more LLMs, both as judges and as evaluated models.
@@ 3 Methodology
As our testbed, we use the TriviaQA dataset (Joshi et al., 2017), consisting of 95K question-answer pairs sourced from 14 trivia and quiz league websites. Each question in the train and validation set is annotated with a list of short answers containing a minimal set of facts and evidence documents collected from Wikipedia and the Web. For our experiments, we use the validation set of the unfiltered partition of the benchmark, using the short answers as reference answers. We use the training set for few-shot examples.
Since experiments require manual annotation of the exam-taker model responses, we use a random sample of $400$ questions from the dataset. In Appendix˜I, we show with a bootstrapping test that this sample size has low variance for our main result. Through experiments described in §˜3, we establish that humans have high agreement on judgements of answers given to the questions in the benchmark.
Figure 2: Difference with human evaluation scores versus alignment metric. The delta evaluation score is the difference between the judge and the human score; y-axes are in log scale. Percent alignment (left) shows a very skewwed distribution, making it difficult to distinguish models. Scott’s $\mathbf{\pi}$ (left) provides a clearer difference between models, and is more indicative of deviation of the gold score.
@@
To understand the strengths and weaknesses of different judges, we consider answers of pre-trained (base) and instruction-tuned (chat) ‘exam-taker models’ across a wide variety of model sizes. In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^1] Accessed via the OpenAI API between Mar 19th, 2024 and Sep 20, 2024. (Achiam et al., 2023) as the exam-taker models. The prompts for the exam-taker models contain five few-shot examples of (question, answer) pairs from the TriviaQA training set. The prompts for the instruction-tuned models additionally include a command signaling the model to answer the given question in a succinct manner similar to the provided examples. The prompts are provided in Appendix˜D.
@@
To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Llama-3.1 (Dubey et al., 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges. To maintain parity with human and judge evaluation, judge prompts were built from human guidelines in Appendix˜G. The judges are instructed to respond with only a single word, ‘‘correct’’ or ‘‘incorrect’’. An overview of all exam-taker models and judge models is shown in Table˜1. For ease of reading, the judge models are depicted in a different font than the exam-taker models.
@@
We use two metrics to quantify alignment between judges: percent agreement and Scott’s Pi coefficient (Scott, 1955). [^2] In an earlier version of this paper, we used Cohen’s kappa (Cohen, 1960) to measure alignment. It has since come to our attention that – despite it’s widespread use – this metric has some well-documented theoretical issues (e.g. Pontius and Millones, 2011; Chicco et al., 2021). For the interested reader, we elaborate on these issues in Appendix B. Percent agreement expresses a simple percentage of the samples on which two annotators agree. Scott’s Pi, denoted as Scott’s $\mathbf{\pi}$, is an alignment metric that corrects for chance agreement between two annotators and is considered to provide a more robust measure of alignment. Details about the computation of both metrics are given in Appendix˜F.
Figure 3: Judge rankings and true/false positives and negatives. (a) Assigned exam-taker model rankings assigned by highly human aligned judges. Contains stays closely to human-assigned rankings, as well as GPT-4 Turbo and Mistral 7B. (b) False positives and negatives across different judge models, in descending order of human alignment. Both false negatives and false positives increase as human alignment decreases, but well-aligned models tend to produce more false positives than false negatives.
@@
As a ground-truth assessment, we obtain human annotations for each exam-taker model answer. The inter-human alignment is calculated between three human judges using the answers to 1200 randomly sampled questions answers; the human guidelines can be found in Appendix˜G. We then determine collective “Human Judgment” through a majority vote.
The average alignment between human evaluators and the majority vote yielded a Scott’s $\mathbf{\pi}$ of $96.2\pm 1.07$, [^3] The coefficient is scaled by $100$ for easier comparison with percentage alignment. while the average percentage agreement was $98.52\%\pm 0.42\%$, exceeding the alignment previously reported in comparable studies (Zeng et al., 2024).
The details of this experiment are mentioned in Appendix˜A. Given this near-perfect alignment score, we consider only one human evaluator per sample for the rest of our experiments, to reduce the overall cost of human annotations. The set of questions for which we obtain human annotations is identical for each exam-taker model.